1. A point charge q = 2.00μC is located at the origin. Find the electric potential due to this charge at point x = 4.0 m.

A. 1500V C. 3500V

B. 2500V D. 4500V


2. Two point charges are arranged along the x-axis; q₁-2.0 μC is at x=0.80m, and q2-2.0 μC at x=0.80m. The net electrical potential measured at the origin due to these point charges is

A. Zero C. 22500V

B. 4500V D. 45000V






1
Expert's answer
2022-04-18T08:50:26-0400

1.

"V=k\frac{q}{r}=9*10^9*\frac{2.00*10^{-6}}{4.0}=4500\:\rm V"

Answer: D

2.

"V=V_1+V_2=k\frac{q_1}{r_1}+k\frac{q_2}{r_2}\\
=9*10^9\left(\frac{2.00*10^{-6}}{0.8}+\frac{2.00*10^{-6}}{0.8}\right)=45000\:\rm V"

Answer: D